1 - 1 Find the eigenvalues of the matrix -41 26 | 136 The eigenvalues are a = bi where Preview Preview b=
Answer #1

14 -4 A = 26 136 She eigenvalues A the esuatin, given by avre 14- 4 26-7 136 (-14-A) 26-)-(-4)(134) = 0 (A+149) (a-24) - 26 1